OpenStudy (anonymous):

120% of what number is 1,331?

hero (hero):

is/of = %/100 1331/x = 120/100 Cross multiply 133100 = 120x 133100/120 = x

OpenStudy (anonymous):

i got 1109.17

OpenStudy (anonymous):

120/100*x = 1331 120x/100 = 1331 Cross multiply 120x = 1331*100 x = 133100/120 x = 1109.17 So, yeah you're right.

OpenStudy (anonymous):

x=1109.166667

OpenStudy (dean.shyy):

Take 1331 and divide it by 120%. Make sure you convert 120% into decimals. That would means you solve the equation 1331/120 = x. Solve for x.